OpenStudy (anonymous):

find all the zeros of f(x) = 3x^3-7x^2+6x-8

OpenStudy (anonymous):

x = 1/6 (1-i sqrt(47)) x = 1/6 (1+i sqrt(47)) x = 2 Would you like an explanation?

OpenStudy (saifoo.khan):

RIght, real root is 2.

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

Factor (x-2) (3 x^2-x+4) x = 2 Then use the quadratic formula.

OpenStudy (anonymous):

for 3x^2-x+4 use quadratic?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I see I see :D thank you.